Main INDEX
Monthly INDEX
PREV
NEXT
User name LeRose
Log entry time 14:35:45 on April 5,2001
Entry number 59977
keyword=Target Down
Upon trying to put beam on the LH2 target in kinematics 24 we discovered the target ladder had moved.
Things looked ok through optics checks and spot on BeO looked ok. After putting in the LH2 target the trouble began.
1st indication was elevated radiation levels in the dump tunnel (ion chamber readings jumped a factor of 10 from before the kinematics change) such that they couldn't give us the current we wanted.
Trying to view beam spot on BeO showed no beam spot. Even with target "out" problems persisted (we must have been hitting the frame).
MCC indicated that all indicators suggested that everything was good right up to the scattering chamber.
We made an access to check the target verniers (there was also some suspicion that there were problems with the OTR) and Dave Meekins and Joe Mitchell concluded that there had been a major shift in the target ladder.
We have gone to restricted access and begun warmup of the target.